# Iker García-Ferrero, Begoña Altuna HiTZ Basque Center for Language Technology - Ixa

### v1

@@
[Uncaptioned image]NoticIA: A Clickbait Article Summarization Dataset in Spanish.
@@ 1 Introduction
In the digital age, the expansion of clickbait headlines has become a pressing concern for both media consumers and publishers. Clickbait refers to sensationalized or misleading headlines designed to lure readers into clicking on a link, often at the expense of accurate reporting and journalistic integrity. Clickbait headlines may mislead readers by exaggerating or distorting the actual content, leading to misinformation and confusion. Clickbait headlines create a need in users to discover specific information, compelling them to click on the article, while the low quality of the piece of news may make them quickly exit the website without engaging with other ads or articles.
@@
The clickbait headlines often lead readers to enter a webpage and scroll down it while they try to find the promised information they are searching for. Such articles often contain a plethora of irrelevant details, with the main idea typically buried at the end. This practice tends to be annoying for the readers and undermines their trust in online news outlets. The ultimate goal of clickbait-led news is to show as many advertisements as possible to the readers in a way of trying to increase the advertisement revenue of the site. For publishers, while clickbait may offer short-term gains in terms of increased web traffic and potential advertising revenue, it carries the risk of damaging their reputation and alienating their reader base. It also negatively impacts advertising revenue for legitimate content creators, who might see their web traffic reduced.
@@
Figure 1: Examples of clickbait headlines. The headline is followed by a long article in which the answer to the headline is located at the end of the article.
@@
The task of summarizing these low-quality articles with clickbait headlines constitutes a great benchmark for Large Language Models (LLMs). Unlike general domain summarization, this task first requires the model to understand the headline and the information it promises. The summary depends on accurately interpreting the headline, as the failure to do so will lead to an incorrect summary. Then, it must successfully navigate through the filler content to identify and extract the key information or idea that is often hidden, not immediately obvious, or, in some cases, completely missing because the headline was misleading. This task demands advanced text understanding and summarization abilities and challenges the models’ capacity to infer and connect diverse pieces of information to fulfil the user’s informational needs as suggested by the clickbait headline. This challenge can be described as an “ultrasummary” generation task, where articles, designed to be as lengthy as possible, are summarized into a single sentence or even a single word. Figure 1 illustrates examples of clickbait headlines from our dataset, together with the human-written summaries.
@@
To the best of our knowledge, [Uncaptioned image]NoticIA is the only clickbait dataset that provides headline-news body-ultra-summary triplets that can be used to assess the performance of information retrieval systems that will help extract crucial information. NoticIA comprises 850 Spanish news articles featuring single-sentence summaries written by humans. Our contributions are the following:
@@ 2 Related works
Most of the efforts aimed at combating clickbaits have focused on clickbait headline detection. Although it is over a decade that the task has attracted the interest of scholars, the Clickbait Challenge[Potthast et al.,2018] in 2017 was a milestone in the advancement of the field. There were 12 teams taking part and the Webis Clickbait Corpus 2017 was released which contained 38,517 annotated tweets and that has been used since then in a list of clickbait detection works such as the work by [Zheng, Yu, and Wu,2021]. In addition, clickbait detection efforts have multiplied in the last lustrum as in the case of [Pujahari and Sisodia,2021], [Liu et al.,2021] or [Wang, Maslim, and Liu,2023].
Our work, however, is more closely related to the clickbait spoiling task. The PAN Clickbait Challenge at SemEval 2023 [Fröbe et al.,2023] consisted in classifying the types of spoilers according to their structure and in generating spoilers according to their type. 23 teams submitted their systems which were tested on the Webis Clickbait Spoiling Corpus 2022, a 5,000 headline-news-spoiler corpus collected from social media.
The task can also be designed in a question-answering manner. In [Heiervang,2022] a title-answering method is proposed. Fine-tuned LLMs are requested to extract an abstractive summary from the news body to answer the question in the clickbait headline. [Johnson et al.,2022] consider both abstractive and extractive summaries to provide an answer for each headline. They fine-tune a RoBERTa model to perform the task. [Intan Maharani, Purwarianti, and Aji,2023], instead, configure the task as a single or multi-span information extraction task for question answering.
@@ 3.1 Corpus collection
Clickbait headlines have proliferated on the web in recent years. However, gathering clickbait headlines is an arduous task. To avoid the need for massive crawling of articles and the subsequent post-processing step to detect articles with clickbait headlines, we utilized two readily available sources of clickbait headlines to build the dataset. The first source is the posting on the X social media platform by the @ahorrandoclick1 user, who manually summarizes articles with clickbait headlines. We gathered 725 articles from this source. Secondly, during the early development of this dataset, we fine-tuned a model to summarize clickbait articles. We used this model to develop a publicly available web app  [^7] https://iker-clickbaitfighter.hf.space/ that users could use to summarize articles given the article’s URL. From the user inputs, we selected an additional 125 articles, for a total of 850 articles in the dataset.
@@ 3.2 Corpus annotation
We employed human annotators to create gold-standard summaries for the dataset. These annotators were PhD students with native-level competence in Spanish. They were instructed to read the headline and identify the question that the clickbait posed for the reader. Subsequently, they were tasked with finding the excerpt of the text that provided the answer to the clickbait headline. The annotators were asked to produce the shortest possible summary that responded solely to the clickbait headline, excluding any other information. Additionally, whenever possible, they were encouraged to directly quote the original text, especially when the summary included direct statements made by individuals. The full guidelines are available in Appendix F. The annotation process took approximately 40 hours. Table 1 lists the average word counts for the headlines, article bodies, and summaries. On average, the article body was summarized with a 98% reduction.
@@ 3.3 Dataset quality assessment
The corpus annotations have been used to measure the quality of the distributed data and guidelines [^8] https://hf.co/datasets/Iker/NoticIA_Human_Validation. A summary of the inter-annotator agreement assessment is presented in Table 2. The overall agreement between the annotators has been high as they have provided the exact same answer in 26% of the cases and have provided answers that partially share information in 48% of the cases. This may imply that, although the news in our dataset tend to present the information intricately, it was easy for humans to find the information the headline refers to.
@@
We have also identified a list of cases in which the annotators have offered different but equally valid answers, which constitutes 18% of the cases. For example (3.3) the annotators have chosen different text strings to refer to the same monetary prize. In (3.3), instead, answers are not directly related.
@@
For what regards the guideline assessment, we venture to say that the guidelines were overall unambiguous and that the request of selecting the minimum amount of words to generate a valid summary will always be interpretable. As we could see from the annotators’ answers, the minimum extent could be understood as the focus of the question in the headline or the minimal well-formed sentence (3.3).
@@
Finally, we identified 8 cases of disagreement. In 3 cases, one of the annotators produced an incorrect summary, likely due to fatigue after annotating multiple examples. In the remaining 5 cases, the disagreement stemmed from contradictory information in the article and different interpretations of this information (3.3). In these instances, determining the correct summary is subject to the reader’s interpretation. In any case, we consider the annotator agreement to be very high, and further experiments in Section 5 demonstrate that there is a very high correlation on model performance when evaluated using the summaries produced by different annotators.
@@ 4.1 Models
Model Family Model name Parameter Count Citation Deepseek deepseek-llm-67b-chat 67B [Bi et al.,2024] Gemma gemma-2b-it 2.51B [Gemma-Team et al.,2024] gemma-7b-it 8.54B [Gemma-Team et al.,2024] Llama2 Llama-2-70b-chat-hf 70B [Touvron et al.,2023] Llama-2-13b-chat-hf 13B [Touvron et al.,2023] tulu-2-dpo-70b 70B [Ivison et al.,2023] Mistral Mistral-7B-Instruct-v0.2 7B [Jiang et al.,2023] Nous-Hermes-2-Mistral-7B-DPO 7B [Teknium,2023] Nous-Hermes-2-Pro-Mistral-7B 7B [Teknium,2023] openchat-3.5-0106 7B [Wang et al.,2023] OpenHermes-2.5-Mistral-7B 7B [Teknium,2023] Solar SOLAR-10.7B-Instruct-v1.0 10.7B [Kim et al.,2023] Nous-Hermes-2-SOLAR-10.7B 10.7B [Teknium,2023] Mixtral Mixtral-8x7B-Instruct-v0.1 46.7B [Jiang et al.,2024] Nous-Hermes-2-Mixtral-8x7B-DPO 46.7B [Teknium,2023] Nous-Hermes-2-Mixtral-8x7B-SFT 46.7B [Teknium,2023] Yi Nous-Hermes-2-Yi-34B 34B [Teknium,2023] Yi-6B-Chat 6B [01.AI et al.,2024] Yi-34B-Chat 34B [01.AI et al.,2024] QWEN Qwen1.5-7B-Chat 7B [Bai et al.,2023] Qwen1.5-14B-Chat 14B [Bai et al.,2023] Qwen1.5-72B-Chat 72B [Bai et al.,2023] ChatGPT gpt-4-0125-preview unknown [OpenAI,2023] gpt-3.5-turbo-0125 unknown [OpenAI,2024]
@@
We evaluate a diverse set of large language models (LLMs), ranging from 2 billion to 70 billion parameters. All the models assessed have been fine-tuned for instruction following. The models include Deepseek [Bi et al.,2024], LLama2 [Touvron et al.,2023], Mistral [Jiang et al.,2023], Solar [Kim et al.,2023], Mixtral [Jiang et al.,2024], StableLM [stability.ai,2023], Yi [01.AI et al.,2024] and Qwen [Bai et al.,2023], as well as further fine-tuned versions of these models which have improved capabilities. The models are listed in Table 4. This selection of models spans the most popular architectures at the time of writing and includes the current best-performing open-source models. It is important to note that most of these models were pretrained with a primary focus on the English or Chinese languages. Our comprehensive evaluation aims to shed light on the performance of the current state-of-the-art LLMs in Spanish. In addition to these open-source models, we also evaluate the GPT3.5 and GPT4 [OpenAI,2023] commercial productions. Although these models are not designed for research, and many details about them remain undisclosed, given their popularity and widespread usage, we consider it valuable to assess their performance on our Spanish ultra-summarization task. We provide further details on the models we use in Appendix D.
@@ 4.2 Task Formulation
We adopt a zero-shot prompt evaluation setup, wherein we construct an instruction that describes the task and the annotation guidelines. The model receives the headline and the article body, and we expect a summary of the article as the output. The prompt used as input for the models is illustrated in Figure 3. The lengthy nature of the inputs, with article bodies averaging 550 words, precludes experimentation with more advanced prompting techniques, such as few-shot learning—this is, providing the model with a few summary examples—or Chain-of-Thought techniques. In the case of models that have been fine-tuned as dialogue systems, these utilize different tokens to represent a conversation, such as using markers like “¡bot¿” and “¡human¿”, or custom system initial prompts. To accommodate these models, we format the input according to the recommendations provided by the authors. The input prompt is set to the user’s input, and we expect the summary to be produced as the assistant’s output. We do not specify any system-defined prompts and, for all the experiments, we use greedy search to generate text.
@@ 4.3 Metrics
As standard in summarization tasks, we use the ROUGE score metric [Lin,2004] to automatically evaluate the summaries produced by the models. ROUGE is a recall-oriented summarization metric that assesses the quality of summarization systems by determining how much of the basic units in the human reference summaries appear in the machine-generated summaries. Our primary metric is ROUGE-1, which considers whole words as the basic units. To compute the ROUGE score, we lowercase both summaries and remove punctuation. In addition to the ROUGE score metric, we also consider the average length of the summaries. For our task, we aim for the summaries to be concise, which is an aspect the ROUGE score does not evaluate. Longer sentences may increase the chance of overlapping words, potentially affecting the ROUGE score. Therefore, we consider both ROUGE-1 and average summary length when evaluating the models. Our goal is to identify a model that achieves the highest possible ROUGE score with the shortest summary length, balancing quality and brevity.
Recent works [Min et al.,2023] have proposed advanced metrics for the automatic evaluation of general-domain summaries. However, our dataset significantly deviates from general domain summarization. Our summaries are highly concise, and we consider any information in the summary that does not contribute towards answering the clickbait headline as incorrect, even if this information appears in the article. Therefore, the set of possible correct summaries is very small, with any accurate summary necessarily including a specific set of words. Thus, we consider the ROUGE score as the best metric for evaluation, along with average sentence length to ensure that the model does not produce overly verbose summaries or include unnecessary information.
@@ 5 Experiments
In this section, we present the evaluation of a broad range of LLMs in both zero-shot and fine-tuning scenarios using the NoticIA dataset. Figure 4 presents a summary of all the results. In this chart, the Y-axis represents the ROUGE score, while the X-axis indicates the average number of words in the summaries produced by each model. As outlined in Section 4.3, our goal is to identify the models that achieve the highest possible ROUGE score with the shortest summary length. Additionally, we establish a human baseline, calculated based on the second set of human annotations detailed in Section 3.3, and compare it against the gold summaries in the dataset. Extended results in numerical format are presented in Appendix E.

### v2

@@
[Uncaptioned image]NoticIA: A Clickbait Article Summarization Dataset in Spanish
@@ 1 Introduction
In the digital age, the expansion of clickbait headlines has become a pressing concern for both media consumers and publishers. Clickbait refers to sensationalized or misleading headlines designed to lure readers into clicking on a link, often at the expense of accurate reporting and journalistic integrity. In clickbait headlines the actual content is exaggerated and distorted, driving readers to misinformation and confusion. Those headlines create a need in users to discover specific information, compelling them to click on the article, although the low quality of the piece of news may make them quickly exit the website without engaging with other ads or articles.
@@
The clickbait headlines often lead readers to enter a webpage and scroll down it while they try to find the promised information they are searching for. Such articles often contain a plethora of irrelevant details, with the main idea typically buried at the end. The ultimate goal of clickbait-led news is to show as many advertisements as possible to the readers in a way of trying to increase the advertisement revenue of the site. This practice tends to be annoying for the readers and undermines their trust in online news outlets. For publishers, while clickbait may offer short-term gains in terms of increased web traffic and potential advertising revenue, it carries the risk of damaging their reputation and alienating their reader base. It also negatively impacts advertising revenue for legitimate content creators, who might see their web traffic reduced.
@@
Figure 1: Examples of clickbait headlines from NoticIA. The headline is followed by a long article in which the answer to the headline is located at the end of the article. We translated two examples into English for illustration.
@@
The task of summarizing these low-quality articles with clickbait headlines constitutes a great benchmark for Large Language Models (LLMs). Unlike general domain summarization, this task first requires the model to understand the headline and the information it promises. The summary depends on accurately interpreting the headline, as the failure to do so will lead to an incorrect summary. Then, it must successfully navigate through the filler content to identify and extract the key information or idea that is often hidden, not immediately obvious, or, in some cases, completely missing because the headline was misleading. This task demands advanced text understanding and summarization abilities and challenges the models’ capacity to infer and connect diverse pieces of information to fulfil the user’s informational needs as suggested by the clickbait headline. This challenge can be described as an “ultrasummary” generation task, where articles, designed to be as lengthy as possible, are summarized into a single sentence or even a single word. Figure 1 illustrates examples of clickbait headlines from our dataset, together with the human-written summaries. We also provide two translated examples to make the task description more accessible.
@@
To the best of our knowledge, [Uncaptioned image]NoticIA is the only clickbait dataset that provides headline-news body-ultrasummary triplets that can be used to assess the performance of information retrieval systems that will help extract crucial information. NoticIA comprises 850 Spanish news articles featuring single-sentence summaries written by humans. Our contributions are the following:
@@ 2 Related works
Most of the efforts aimed at combating clickbaits have focused on clickbait headline detection. Although it is over a decade that the task has attracted the interest of scholars, the Clickbait Challenge ([Potthast et al., 2018] in 2017 was a milestone in the advancement of the field. There were 12 teams taking part and the Webis Clickbait Corpus 2017 was released, which contained 38,517 annotated tweets and that has been used since in a list of clickbait detection works such as the work by [Zheng, Yu, and Wu, 2021]. In addition, clickbait detection efforts have multiplied in the last lustrum as in the case of [Pujahari and Sisodia, 2021], [Liu et al., 2021] or [Wang, Maslim, and Liu, 2023].
Our work, however, is more closely related to the clickbait spoiling task. The PAN Clickbait Challenge at SemEval 2023 ([Fröbe et al., 2023] consisted in classifying the types of spoilers according to their structure and in generating spoilers according to their type. 23 teams submitted their systems, which were tested on the Webis Clickbait Spoiling Corpus 2022, a 5,000 headline-news-spoiler corpus collected from social media.
The task can also be designed in a question-answering manner. In [Heiervang, 2022] a title-answering method is proposed. Fine-tuned LLMs are requested to extract an abstractive summary from the news body to answer the question in the clickbait headline. [Kurenkov et al., 2022] consider both abstractive and extractive summaries to provide an answer for each headline. They fine-tune a RoBERTa model to perform the task. [Intan Maharani, Purwarianti, and Aji, 2023], instead, configure the task as a single or multi-span information extraction task for question answering.
@@ 3.1 Corpus collection
Clickbait headlines have proliferated on the web in recent years. However, gathering clickbait headlines is an arduous task. To avoid the need for massive crawling of articles and the subsequent post-processing step to detect articles with clickbait headlines, we utilized two readily available sources of clickbait headlines to build the dataset. The first source is the posting on the X social media platform by the @ahorrandoclick1 user, who manually summarizes articles with clickbait headlines, from which we gathered 725 articles. Secondly, during the early development of this dataset, we fine-tuned a model to summarize clickbait articles. We used this model to develop a publicly available web app [^7] https://iker-clickbaitfighter.hf.space/ that users could use to summarize articles given the article’s URL. From the user inputs, we selected an additional set of 125 articles, for a total of 850 articles in the dataset.
@@ 3.2 Corpus annotation
We employed human annotators to create gold standard summaries for the dataset. These annotators were PhD students with native-level competence in Spanish. They were instructed to read the headline and identify the question that the clickbait posed for the reader. Subsequently, they were tasked with finding the excerpt of the text that provided the answer to the clickbait headline. The annotators were asked to produce the shortest possible summary that responded solely to the clickbait headline, excluding any other information. Additionally, whenever possible, they were encouraged to directly quote the original text, especially when the summary included direct statements made by individuals. The full guidelines are available in Appendix F. The annotation process took approximately 40 hours. Table 1 lists the average word counts for the headlines, article bodies, and summaries. On average, the article body was summarized with a 98% reduction.
@@ 3.3 Dataset quality assessment
The corpus annotations have been used to measure the quality of the distributed data and guidelines. [^8] https://hf.co/datasets/Iker/NoticIA_Human_Validation The test partition of our dataset was annotated by a second annotator to assess whether humans can provide aligned answers based on the guidelines. A summary of the inter-annotator agreement assessment is presented in Table 2.
The overall agreement between the annotators has been high as they have provided the exact same answer in 26% of the cases and have provided answers that partially share information in 48% of the cases. This may imply that, although the news in our dataset tend to present the information intricately, it was easy for humans to find the information the headline refers to.
@@
We have also identified a list of cases in which the annotators have offered different but equally valid answers, which constitutes the 18% of the cases. For example (3.3) the annotators have chosen different text strings to refer to the same monetary prize. In (3.3), instead, answers are not directly related as one refers to cold air and the other to unusually low temperatures.
@@
For what regards the guideline assessment, we venture to say that the guidelines were overall unambiguous and that the request of selecting the minimum amount of words to generate a valid summary will always be interpretable. As we could see from the annotators’ answers, the minimum extent could be understood as the minimal well-formed sentence or the focus of the question in the headline (3.3).
@@
Finally, we identified 8 cases of disagreement. In 3 cases, one of the annotators produced an incorrect summary, likely due to fatigue after annotating multiple examples. In the remaining 5 cases, the disagreement stemmed from contradictory information in the article and different interpretations of this information (3.3). In these instances, determining the correct summary is subject to the reader’s interpretation. In any case, we consider the annotator agreement to be very high, and further experiments in Section 5.2.1 demonstrate that there is a very high correlation on model performance when evaluated using the summaries produced by different annotators.
@@ 4.1 Models
Model Family Model name Parameter Count Citation Deepseek deepseek-llm-67b-chat 67B ([Bi et al., 2024] Gemma gemma-2b-it 2.51B ([Gemma-Team et al., 2024] gemma-7b-it 8.54B ([Gemma-Team et al., 2024] Llama2 Llama-2-70b-chat-hf 70B ([Touvron et al., 2023] Llama-2-13b-chat-hf 13B ([Touvron et al., 2023] tulu-2-dpo-70b 70B ([Ivison et al., 2023] Llama3 Meta-Llama-3-8B-Instruct 70B ([AI@Meta, 2024] Meta-Llama-3-8B-Instruct 8B ([AI@Meta, 2024] Mistral Mistral-7B-Instruct-v0.2 7B ([Jiang et al., 2023] Nous-Hermes-2-Mistral-7B-DPO 7B ([Teknium, 2023] Nous-Hermes-2-Pro-Mistral-7B 7B ([Teknium, 2023] openchat-3.5-0106 7B ([Wang et al., 2023] OpenHermes-2.5-Mistral-7B 7B ([Teknium, 2023] Solar SOLAR-10.7B-Instruct-v1.0 10.7B ([Kim et al., 2023] Nous-Hermes-2-SOLAR-10.7B 10.7B ([Teknium, 2023] Mixtral Mixtral-8x7B-Instruct-v0.1 46.7B ([Jiang et al., 2024] Nous-Hermes-2-Mixtral-8x7B-DPO 46.7B ([Teknium, 2023] Nous-Hermes-2-Mixtral-8x7B-SFT 46.7B ([Teknium, 2023] Yi Nous-Hermes-2-Yi-34B 34B ([Teknium, 2023] Yi-6B-Chat 6B ([01.AI et al., 2024] Yi-34B-Chat 34B ([01.AI et al., 2024] QWEN Qwen1.5-7B-Chat 7B ([Bai et al., 2023] Qwen1.5-14B-Chat 14B ([Bai et al., 2023] Qwen1.5-72B-Chat 72B ([Bai et al., 2023] ChatGPT gpt-4o-2024-05-13 unknown ([OpenAI, 2024b] gpt-4-turbo-2024-04-09 unknown ([OpenAI, 2023] gpt-3.5-turbo-0125 unknown ([OpenAI, 2024a]
@@
We evaluate a diverse set of large language models (LLMs), ranging from 2 billion to 70 billion parameters. All the models assessed have been fine-tuned for instruction following. The models include Deepseek ([Bi et al., 2024], LLama2 ([Touvron et al., 2023], Llama3 ([AI@Meta, 2024], Mistral ([Jiang et al., 2023], Solar ([Kim et al., 2023], Mixtral ([Jiang et al., 2024], StableLM ([stability.ai, 2023], Yi ([01.AI et al., 2024] and Qwen ([Bai et al., 2023], as well as further fine-tuned versions of these models which have improved capabilities. The models are listed in Table 4. This selection of models spans the most popular architectures at the time of writing and includes the current best-performing open-source models. Yet, it is important to note that most of these models were pretrained with a primary focus on the English or Chinese languages. Our comprehensive evaluation aims to shed light on the performance of the current state-of-the-art LLMs in Spanish. In addition to these open-source models, we also evaluate the GPT-3.5, GPT-4 and GPT-4o commercial products. Although these models are not designed for research, and many details about them remain undisclosed, given their popularity and widespread usage, we consider it valuable to assess their performance on our Spanish ultra-summarization task. We provide further details on the models we use in Appendix D.
@@ 4.2 Task Formulation
We adopt a zero-shot prompt evaluation setup, wherein we construct an instruction that describes the task and the annotation guidelines. The model receives the headline and the article body, and we expect a summary of the article as the output. The prompt used as input for the models is illustrated in Figure 3, a linguist expert on the summarization task crafted this prompt. The lengthy nature of the inputs, with article bodies averaging 550 words, precludes experimentation with more advanced prompting techniques, such as few-shot learning—this is, providing the model with a few summary examples—or Chain-of-Thought techniques. In the case of models that have been fine-tuned as dialogue systems, these utilize different tokens to represent a conversation, such as using markers like “¡bot¿” and “¡human¿”, or custom system initial prompts. To accommodate these models, we format the input according to the recommendations provided by the authors. The input prompt is set to the user’s input, and we expect the summary to be produced as the assistant’s output. We do not specify any system-defined prompts and, for all the experiments, we use greedy search to generate text.
@@ 4.3 Metrics
As standard in summarization tasks, we use the ROUGE score metric ([Lin, 2004] to automatically evaluate the summaries produced by the models. ROUGE is a recall-oriented summarization metric that assesses the quality of summarization systems by determining how much of the basic units in the reference summaries [^9] First annotator’s summaries in this work. appear in the machine-generated summaries. Our primary metric is ROUGE-1, which considers whole words as the basic units. To compute the ROUGE score, we lowercase both summaries and remove punctuation. In addition to the ROUGE score metric, we also consider the average length of the summaries. For our task, we aim for the summaries to be concise, which is an aspect the ROUGE score does not evaluate. Longer sentences may increase the chance of overlapping words, potentially affecting the ROUGE score. Therefore, we consider both ROUGE-1 and average summary length when evaluating the models. Our goal is to identify a model that achieves the highest possible ROUGE score with the shortest summary length, balancing quality and brevity.
Recent works ([Min et al., 2023] have proposed advanced metrics for the automatic evaluation of general-domain summaries. However, our dataset significantly deviates from general domain summarization. Our summaries are highly concise, and we consider any information in the summary that does not contribute towards answering the clickbait headline as incorrect, even if this information appears in the article. Therefore, the set of possible correct summaries is very small, with any accurate summary necessarily including a specific set of words. Thus, we consider the ROUGE score as the best metric for evaluation, along with average sentence length to ensure that the model does not produce overly verbose summaries or include unnecessary information.
@@ 5 Experiments
In this section, we present the evaluation of a broad range of LLMs in both zero-shot and fine-tuning scenarios using the NoticIA dataset. Figure 4 presents a summary of all the results, both in the zero-shot and fine-tuning (ClickbaitFighter models) settings. In this chart, the Y-axis represents the ROUGE score, while the X-axis indicates the average number of words in the summaries produced by each model. As outlined in Section 4.3, our goal is to identify the models that achieve the highest possible ROUGE score with the shortest summary length. Additionally, we establish a human baseline, calculated on the agreement between the human annotators detailed in Section 3.3, and compare it against the gold summaries in the dataset. Extended results in numerical format are presented in Appendix E.